Snow Cone Inc. has determined its cost of each source of capital and optimal capital structure as followed.

Source of Capital

Proportions

Before-tax Cost

Long-term Debt

$45,000

8%

Common stock

$55,000

16%

The firm has 21% tax rate. How much is their weighted average cost of capital?

15.28%

12.40%

11.64 %

13.00%

S'more Fest Enterprises has the following project that it can invest $30,000 today. The project will generate $9,000, $10,000; $8,000; $12,000; and $13,000 in 5 years respectively. What is their payback period?

Question options:

4.12 years

3.25 years

3.0 years

5.06 years
Answer #1

A)Correct option is "C"- 11.64%

Cost Proportion weights Cost *weight
After tax cost of Debt

8(1-.21) = 6.32%

45000 45000/100000=.45 6.32*.45=2.84
common stock 16% 55000 55000/100000=.55

16*.55= 8.8

100000 11.64%

B)Correct option is "B"- 3.25 years

Year Cash Flow Cumulative cash flow
0 -30000 -30000
1 9000 -30000+9000=-21000
2 10000 -21000+10000=-11000
3 8000 -11000+8000=-3000
4 12000 -3000+12000=9000
5 13000 9000+13000=22000

Payback period =year up to which cumulative cash flow is negative +(cumulative cash flow of that period /cash flow of next year)

= 3+ (3000/12000)

= 3 + .25

   = 3.25 years